### v3

# Study of hybrid stars with nonstrange quark matter cores

Abstract
In this work, under the hypothesis that quark matter may not be strange (Holdom et al., 2018), we adopt a modification of the coupling constant of the four-quark scalar interaction $G\rightarrow G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$ in the 2-flavor Nambu-Jona-Lasinio (NJL) model to study nonstrange hybrid stars, where $G_{1}$ and $G_{2}$ are two parameters constrained by using the lattice QCD simulation results at the critical temperature and zero chemical potential. The Maxwell construction is used to describe the first-order confinement-deconfinement phase transition in hybrid stars. With recent measurements on neutron star mass, radius, and tidal deformability, the hybrid equation of states are constrained. It is found that pure nonstrange quark matter cores can exist in hybrid stars, possessing $0.026-0.04$ solar mass. The maximum hybrid star mass in the framework of the modified NJL model is about 0.1 solar mass lighter than that in the conventional 2-flavor NJL model. It is argued that the binary neutron stars in GW170817 should be hadron stars.
thanks: E-mail: licm@zzu.edu.cnthanks: E-mail: zhaoyapeng2013@hotmail.com
1 Introduction
The binary neutron star (BNS) merger GW170817 opened a new era of multi-messenger astronomy  (Abbott & et al., 2017a, b; Margalit & Metzger, 2017; Bauswein et al., 2017; Shibata et al., 2017; Annala et al., 2018; Fattoyev et al., 2018; Paschalidis et al., 2018; Zhou et al., 2018; Ruiz et al., 2018; Radice et al., 2018; Rezzolla et al., 2018; Nandi & Char, 2018; Zhu et al., 2018; Ai et al., 2018; Ma et al., 2019; Zhang, 2020; Li et al., 2020; Miao et al., 2021; Tan et al., 2022; Zou & Huang, 2022; Li et al., 2022b; Zhang et al., 2023). More and more astronomical observations on neutron stars arise, facilitating the study of neutron star structure and equation of state (EOS). As natural laboratories to investigate the dense strongly interacting matter, neutron stars have been attracting much attention in astrophysics and theoretical physics. In general, the characteristic temperature of neutron stars can be well described by zero temperature approximation, due to their excessively high energy density in the interior, thus the quantum chromodynamics (QCD) needs to be employed to study the EOS in neutron stars. It is believed that the density in the core of neutron stars could reach 5-10 $\rho_{0}$, where $\rho_{0}=0.16$ fm-3 is the nuclear saturation density (Lattimer & Prakash, 2004; Lattimer, 2021). As a result, the hadron-quark phase transition is very likely to happen and the deconfined quark matter will appear. In this case, neutron stars are essentially hybrid stars. However, it is difficult to give a unified description of the hadronic matter, quark matter and the hadron-quark phase transition with a single theoretical framework. Thus the hadronic matter and quark matter in hybrid stars are separately described with different EOSs at present, and a certain construction scheme needs to be employed to combine them to get a complete EOS.
As we know, the results of different effective models can be quantitatively or even qualitatively different. Even for the same model, if different modifications are taken into account, the results can also be different. For example, Fig. 10 of Özel & Freire (2016) shows that the EOSs given by different effective models are different from each other, and the corresponding mass-radius ($M-R$) relations of neutron stars are also different. Thus there is not a definite answer to the EOS of dense strongly interacting matter at zero temperature at present.
To describe the hadronic matter in hybrid stars, the EOS developed by Akmal, Pandharipande & Ravenhall (APR) with $A18+\delta\nu+UIX^{\ast}$ interaction is employed in this work (Akmal et al., 1998), in which the Argonne $\nu_{18}$ two-nucleon interaction and boost corrections to the two-nucleon interaction as well as the three-nucleon interaction are taken into account. The hadronic matter in the context of the APR model is one kind of charge-neutral and beta-stable fluid whose pressure and baryon chemical potential are equilibrated. However, to describe the quark matter in hybrid stars, the lattice QCD is confronted with difficulties at low-temperature and high-density regions because of the “sign problem”, thus we need to use effective models, such as the Nambu-Jona-Lasinio (NJL) model  (Klevansky, 1992; Buballa, 2005; Li et al., 2019b; Liu et al., 2021; Zhang et al., 2022; Huang & Zhuang, 2023), which manifests the spontaneous breaking of chiral symmetry.
In the framework of the NJL-type model, many studies focused on hybrid stars (Ayriyan et al., 2021; Blaschke et al., 2020; Alvarez-Castillo et al., 2016; Pfaff et al., 2022), aiming to explain the observed two-solar-mass (2 $M_{\odot}$) compact stars. In Ayriyan et al. (2021); Blaschke et al. (2020); Alvarez-Castillo et al. (2016), the 2-flavor NJL-type model was used to consider the scalar quark-antiquark interaction, anti-triplet scalar diquark interactions and vector quark-antiquark interactions  (Ayriyan et al., 2021), a chemical potential dependence of the vector mean-field coupling $\eta(\mu)$ and a chemical potential-dependent bag constant $B(\mu)$  (Blaschke et al., 2020), multiquark (4- and 8-quark) interactions  (Alvarez-Castillo et al., 2016), respectively. In Pfaff et al. (2022), the 3-flavor SU(3) NJL model was adopted with the four-quark scalar, vector-isoscalar and vector-isovector interactions as well as the ’t Hooft interaction. Different from the above studies, we adopt a modification of the coupling constant of four-quark scalar interactions as $G\rightarrow G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$, which can be regarded as a representation of an effective gluon propagator (See Sec. 2 for specific analysis).
As for the hadron-quark phase transition, the most widely used approach is the Maxwell construction  (Endo et al., 2006; Hempel et al., 2009; Yasutake et al., 2009), assuming that the first-order phase transition occurs (Glendenning, 1992; Bhattacharyya et al., 2010) and stable quark matter cores exist in hybrid stars. However, many studies showed that hybrid stars are unstable against oscillations in this case, because star masses decrease with the increase of the central density, thus quark matter cores may not exist in neutron-star interiors  (Özel, 2006; Hoyos et al., 2016; Qin et al., 2023). In Özel (2006) and Hoyos et al. (2016), the theoretical modeling of bursting neutron-star spectra and top-down holographic model for strongly interacting quark matter were employed, respectively, to demonstrate that the 2 $M_{\odot}$ neutron star has ruled out soft EOSs of neutron-star matter, and no quark matter exists in massive neutron stars. Recently, it has been argued that as the density increases, the boundaries of hadrons disappear gradually and the corresponding phase transition is a crossover  (Baym et al., 2018). According to this assumption, the three-window modeling (Masuda et al., 2013a, b) in the crossover region was proposed. Many studies has constructed hybrid EOSs in this scheme, and the corresponding maximum masses of hybrid stars are compatible with 2 $M_{\odot}$  (Kojo et al., 2015; Li et al., 2017, 2018b, 2018a, 2022a; Qin et al., 2023).
In addition to theoretical studies of hadron-quark phase transitions and hybrid EOSs, astronomical observations of neutron star masses, radii, and tidal deformability have also placed constraints on numerous EOSs. Some massive neutron star observations such as PSR J0348+0432 (Antoniadis et al., 2013) and PSR J0740+6620 (Cromartie et al., 2020) require EOSs should not be too soft, but the tidal deformability constrained in BNS merger event GW170817 indicates the EOSs should not be too stiff (Abbott & et al., 2017a, 2018). Recently, the joint $M-R$ observations of neutron stars from NASA’s Neutron Star Interior Composition Explorer (NICER) missions have also imposed some constraints on these EOSs  (Riley et al., 2019; Miller et al., 2019b; Riley et al., 2021; Miller et al., 2021). In Bauswein et al. (2019); Miao et al. (2020), the authors claim that the gravitational-wave (GW) emission of GW170817 supports a first-order hadron-quark phase transition at supranuclear densities.
In this work, inspired by a recent work that the quark matter may not be strange (Holdom et al., 2018), we will study nonstrange hybrid EOSs and hybrid stars with the Maxwell construction. The hadronic EOS and quark EOS are described by the APR model and a modified 2-flavor NJL model, respectively. The parameter space of $G_{1}$ and $G_{2}$ in the modified NJL model will be fixed according to the lattice results at zero chemical potential  (Laermann & Philipsen, 2003). With recent measurements on neutron star mass, radii, and tidal deformability, the hybrid EOSs will be constrained to get the parameter space of the current quark mass. To ensure that hybrid stars are stable against oscillations, maximum masses of hybrid stars and the masses of their quark matter cores are determined.
It is known that the Bayesian analysis is a good approach to constrain the EOSs. Researchers have obtained important information about the EOS of QCD in this way (Ayriyan et al., 2021; Blaschke et al., 2020; Alvarez-Castillo et al., 2016; Pfaff et al., 2022; Alvarez-Castillo et al., 2020; Ayriyan et al., 2019, 2015; Miller et al., 2019a). However, considering that the lagrangian of the NJL model is convenient for numerical calculation, in this work we have performed calculations focusing on the EOS to get the corresponding hybrid star $M-R$ relations and tidal deformability, and then compared them with the relevant neutron star astronomical observations. The model parameters and properties of hybrid stars are constrained as well.
This paper is organized as follows. In Sec.2, the modified NJL model for nonstrange quark matter is briefly introduced, and the EOSs of quark matter are derived. In Sec.3, the Maxwell construction is used to get hybrid EOSs. With recent astronomical observations of neutron star mass, radius, and tidal deformability, we constrain the hybrid EOSs. For comparison, the $M-R$ relations and tidal deformability results of hybrid stars from six representative hybrid EOSs are presented. Finally, a brief summary is given in Sec.4.
2 EOS of nonstrange quark matter
As an effective model to describe cold dense quark matter, the NJL model (Klevansky, 1992; Buballa, 2005) is widely used in the study of hybrid stars and quark stars. In this work, we consider a modified version, in which the Lagrangian has the following form:
[…]
where $m$ is the current quark mass (because of an exact isospin symmetry between $u$ and $d$ quarks adopted in this work, $m_{\rm u}=m_{\rm d}=m$). Different from the normal NJL model, we adopt $G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$ to represent the four-fermion coupling strength, where $\langle\bar{\psi}\psi\rangle$ is the quark condensate. The term $(G_{1}+G_{2}\langle\bar{\psi}\psi\rangle)[(\bar{\psi}\psi)^{2}+(\bar{\psi}i\gamma^{5}\tau\psi)^{2}]$ describes interactions in scalar-isoscalar and pseudoscalar-isovector channels.
In the following we will clarify the modification in detail. Based on our current knowledge of strong interactions, the coupling constant $G$ in the normal NJL model can be regarded as a representation of an effective gluon propagator. In light of QCD theory, the quark and gluon propagators should satisfy their respective Dyson-Schwinger (DS) equations, and these two equations are coupled with each other. It is demonstrated that quark propagators in the Nambu phase and Wigner phase are very different from each other  (Cui et al., 2018; Xu et al., 2018; Li et al., 2019a), so it can be inferred that the corresponding gluon propagators in these two phases are also different (Hong-Shi & Wei-Min, 2006). However, in the normal NJL model, $G$ is simplified as a constant, remaining the same in these two phases. In addition, according to simulations of lattice QCD, the gluon propagator should vary with temperature, although its dependence on the chemical potential is still uncertain. In the normal NJL model, as a representation of an effective gluon propagator, the coupling constant $G$ is “static”, and thus cannot fulfill the requirement of lattice QCD.
In the QCD sum rule approach (Reinders et al., 1985), it is argued that the full Green function can be divided into two parts: the perturbative part and nonperturbative part. The condensates can be expressed as various moments of nonperturbative Green function. As a result, the most general form of the “nonperturbative” gluon propagator is
$D_{\mu\nu}^{\rm{npert}}\equiv D_{\mu\nu}^{\rm{full}}-D_{\mu\nu}^{\rm{pert}}\equiv c_{1}\langle\bar{\psi}\psi\rangle+c_{2}\langle G^{\mu\nu}G_{\mu\nu}\rangle+...,$ (2)
where $\langle G^{\mu\nu}G_{\mu\nu}\rangle$ refers to the gluon condensate, $c_{1}$ and $c_{2}$ are coefficients which can be calculated in the QCD sum rule approach  (Steele, 1989; Pascual & Tarrach, 1984), and the ellipsis represents the contributions from other condensates, such as the mixed quark-gluon condensate. Among these condensates, the quark condensate possesses the lowest dimension, and a nonzero value of it, in the chiral limit, precisely signifies the dynamical chiral symmetry breaking. Therefore, it plays the most important role in the QCD sum rule approach. In this work, we will deal with its contribution separately, and the contribution of other condensates is simplified into the perturbative part of the gluon propagator. In the normal NJL model, it is equivalent to a modification of the coupling constant $G$ in the following way  (Jiang et al., 2012; Cui et al., 2013, 2014b; Shi et al., 2016; Wang et al., 2016; Fan et al., 2017, 2019; Li et al., 2018b),
$G\rightarrow G_{1}+G_{2}\langle\bar{\psi}\psi\rangle.$ (3)
Now the coupling strength $G$ will depend on both $u$ and $d$ quark condensates via this modification, where $G_{2}$ refers to the weight factor of the influence of the quark propagator on the gluon propagator. It should be noted that following this modification, we need to correspondingly change the term $G\langle\bar{\psi}\psi\rangle^{2}$ in the mean-field thermodynamic potential of the normal NJL model to ensure the consistency in thermodynamics. Similar to Gorenstein & Yang (1995), the term $G\langle\bar{\psi}\psi\rangle^{2}$ is replaced by the integral over $\langle\bar{\psi}\psi\rangle$ of $2\langle\bar{\psi}\psi\rangle{\rm d}(G\langle\bar{\psi}\psi\rangle)/{\rm d}\langle\bar{\psi}\psi\rangle$, in which d…/d… denotes derivative, and the result is $G_{1}\langle\bar{\psi}\psi\rangle^{2}+4G_{2}\langle\bar{\psi}\psi\rangle^{3}/3$. In this way, the mean-field thermodynamic potential has the following form,
$\Omega(T,\{\mu\},\{\langle\bar{\psi}\psi\rangle\})=$ $\Omega_{\rm M}(T,\mu)+G_{1}\langle\bar{\psi}\psi\rangle^{2}+\frac{4}{3}G_{2}\langle\bar{\psi}\psi\rangle^{3}$ (4)
$+{\rm const},$
where $\Omega_{\rm M}$ denotes the contribution of a gas of quasiparticles,
$\Omega_{\rm M}=$ $-2N_{\rm c}N_{\rm f}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}\{T\,{\rm ln}(1+{\rm exp}(-\frac{1}{T}(E_{p}-\mu)))$ (5)
$+T\,{\rm ln}(1+{\rm exp}(-\frac{1}{T}(E_{p}+\mu)))+E_{p}\},$
where $E_{p}=\sqrt{\overrightarrow{p}^{2}+M^{2}}$ is the quark on-shell energy. Note that when $G={\rm const}$ (i.e., $G_{1}=G$ and $G_{2}=0$), our model reproduces the condensate term in the normal NJL model.
The effective mass of the constituent quark is now given by
$M$ $=$ $m-2(G_{1}+G_{2}\langle\bar{\psi}\psi\rangle)\langle\bar{\psi}\psi\rangle.$ (6)
The quark condensate $\langle\bar{\psi}\psi\rangle$ and the particle number density $\rho$ can be derived from $\Omega$ as
$\langle\bar{\psi}\psi\rangle$ $=$ $\frac{\partial\Omega}{\partial m}$
$=$ $-2N_{\rm c}N_{\rm f}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}\frac{M}{E_{p}}[1-n_{p}(T,\mu)-\overline{n}_{p}(T,\mu)],$
$\rho$ $=$ $-\frac{\partial\Omega}{\partial\mu}$ (8)
$=$ $2N_{\rm c}N_{\rm f}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}(n_{p}(T,\mu)-\overline{n}_{p}(T,\mu)),$
where $n_{p}(T,\mu)$ and $\overline{n}_{p}(T,\mu)$ are the Fermi occupation numbers of quarks and antiquarks, respectively, which are defined as
$n_{p}(T,\mu)$ $=$ $[{\rm exp}^{(E_{p}-\mu)/T}+1]^{-1},$ (9)
$\overline{n}_{p}(T,\mu)$ $=$ $[{\rm exp}^{(E_{p}+\mu)/T}+1]^{-1}.$ (10)
Because the NJL model cannot be renormalized, the proper-time regularization is adopted in the following calculations. In addition, we need to fix the parameter set ($\Lambda_{\rm UV}$, $G$) to fit experimental data ($f_{\pi}=92$ MeV, $M_{\pi}=135$ MeV) at zero temperature and chemical potential. The parameter fixing process is similar to that of Cui et al. (2014a).
Although the lattice QCD is confronted with the “sign problem” at finite chemical potentials, the simulating results at zero chemical potential can still help us determine the values of $G_{1}$ and $G_{2}$. According to the simulations of lattice QCD, the chiral phase transition at zero chemical potential is a crossover, and the corresponding pseudo-critical point is located at $T_{\rm pc}=173\pm 8$ MeV in the 2-flavor case  (Laermann & Philipsen, 2003). Different from the meaning of the so-called “critical point” in the case of first-order phase transition, the “pseudo-critical point” here refers to the condition that the crossover occurs, and its position can be identified by the peak of susceptibilities, such as the chiral susceptibility $\chi_{s}$ in Laermann & Philipsen (2003), which is defined as $\chi_{s}=-\partial\langle\bar{\psi}\psi\rangle/\partial m$  (Du et al., 2013).
Figure 1: The chiral susceptibilities $\chi_{\rm s}$ versus temperature for different $G_{1}$. $m$ is fixed as 7.5 MeV. The three vertical lines denote the peak place of $\chi_{\rm s}$, i.e., $T_{\rm pc}=165,181,247.1$ MeV.
Figure 2: The scaled quark condensates versus $T$ for different $G_{1}$. $m$ is fixed as 7.5 MeV. The corresponding pseudo-critical points are also marked on these curves, respectively.
According to our calculations, when the current quark mass $m$ is in the range of 3.5 – 10 MeV, $T_{\rm pc}$ will be larger than 181 MeV in the conventional NJL model. Generally, $T_{\rm pc}$ will decrease as $G_{1}$ decreases. In Fig. 1, the chiral susceptibility for different $G_{1}$ is plotted, taking $\mu=0$ and $m=7.5$ MeV. As $G_{1}$ varies from 4.358 to 4.552 GeV-2, $T_{\rm pc}$ changes from 165 to 181 MeV, satisfying the constraint from lattice simulations. However, when $G_{1}=G=5.29$ GeV-2, $T_{\rm pc}$ in the conventional NJL model is 247.1 MeV, which is much larger than the corresponding value of lattice simulations.
The QCD sum rule at the renormalization scale of 1 GeV suggests the $u$ quark condensate should be $229\pm 33$ MeV3  (Dosch & Narison, 1998) or $242\pm 15$ MeV3  (Jamin, 2002). Considering this constraint, the whole parameter sets of the modified 2-flavor NJL model in this work are shown in Table 1. For a certain current quark mass, we take two boundary values for $G_{1}$, corresponding to the cases with $T_{\rm pc}=165$ and 181 MeV, respectively. Similar to previous studies under the proper-time regularization (Klevansky, 1992; Kohyama et al., 2015; Li et al., 2017), the effective quark masses in Table 1 are in a range of 190 – 250 MeV, which are smaller than 300 MeV. As a comparison, it has been shown that for the same current quark mass, the three-momentum cutoff regularization leads to a larger effective quark mass and a smaller momentum cutoff  (Klevansky, 1992; Ratti et al., 2007; Kohyama et al., 2015, 2016).
Table 1: Parameter sets in this work.
$\quad m$ $\,\,\Lambda_{\rm UV}$ $\,\,M$ ${-}\langle\bar{u}u\rangle^{\frac{1}{3}}$ $\quad\,\,G$ $\quad\,\,\,G_{1}$ $\quad\,\,\,G_{2}$
$[{\rm MeV}]$ $[{\rm MeV}]$ $[{\rm MeV}]$ $\,\,[{\rm MeV}]$ $[{\rm GeV}^{-2}]$ $[{\rm GeV}^{-2}]$ $[{\rm GeV}^{-5}]$
    4.5  1151    191191\,\,\,191191      258258\,\,\,\,\,258258    2.7272.727\,\,\,2.7272.727 2.5632.563$\quad$ 2.5632.563 $\,\,\,-4.791$
    4.5  1151    191191\,\,\,191191      258258\,\,\,\,\,258258    2.7272.727\,\,\,2.7272.727 2.6272.627$\quad$ 2.6272.627 $\,\,\,-2.924$
    5.8  1002    206206\,\,\,206206      237237\,\,\,\,\,237237    3.7613.761\,\,\,3.7613.761 3.3753.375$\quad$ 3.3753.375 $\,\,\,-14.50$
    5.8  1002    206206\,\,\,206206      237237\,\,\,\,\,237237    3.7613.761\,\,\,3.7613.761 3.4853.485$\quad$ 3.4853.485 $\,\,\,-10.36$
    7.5  873    225225\,\,\,225225      217217\,\,\,\,\,217217    5.2905.290\,\,\,5.2905.290 4.3584.358$\quad$ 4.3584.358 $\,\,\,-45.29$
    7.5  873    225225\,\,\,225225      217217\,\,\,\,\,217217    5.2905.290\,\,\,5.2905.290 4.5524.552$\quad$ 4.5524.552 $\,\,\,-35.86$
    9.5  774    250250\,\,\,250250      201201\,\,\,\,\,201201    7.3997.399\,\,\,7.3997.399 5.2265.226$\quad$ 5.2265.226 $\,\,\,-133.8$
    9.5  774    250250\,\,\,250250      201201\,\,\,\,\,201201    7.3997.399\,\,\,7.3997.399 5.5725.572$\quad$ 5.5725.572 $\,\,\,-112.5$
Fig. 2 plots the scaled order parameter of chiral phase transition ($\langle\bar{\psi}\psi\rangle/\langle\bar{\psi}\psi\rangle_{0}$) versus temperature when $m=7.5$ MeV. We can find that $\langle\bar{\psi}\psi\rangle/\langle\bar{\psi}\psi\rangle_{0}$ decreases smoothly from one to zero as temperature increases, thus the transition at $\mu=0$ is the crossover, consistent with the simulation result of lattice QCD.
Now we extend our calculation to finite chemical potentials at $T=0$ to get EOSs of the quark matter. After solving Eq. (6) with the modification of Eq. (3), we can get the dependence of effective quark mass $M$ on the chemical potential, which is shown in Fig. 3. It can be seen that for the dense quark systems with a fixed $T_{\rm pc}$, a larger $m$ leads to a larger effective quark mass in the vacuum, and thus a larger gap will emerge when the chiral phase transition occurs. Specifically, when $T_{\rm pc}=181$ MeV, the crossover occurs for the systems with $m<9$ MeV, and the first order phase transition occurs for the systems with $m>9$ MeV. The critical chemical potential ($\mu_{\rm c}$) is around 285 MeV. As a comparison, we also present the results for a fixed $m$ but with different values of $G_{1}$ in Fig. 3. As $G_{1}$ increases, $\mu_{\rm c}$ will also increase. Note that the green dashed line in Fig. 3 corresponds to normal NJL model, in which the pseudo-critical chemical potential is about 350 MeV. It is much larger than the corresponding results of those cases with $G_{1}\in(4.430,4.552)$ GeV-2.
Figure 3: The effective mass of quarks versus chemical potentials $\mu$ at $T=0$. Note that m58, m75, m95 correspond to three cases with $m=5.8$, 7.5, 9.5 MeV, respectively, in which we have taken $T_{\rm pc}=181$ MeV. Similarly, in the three cases of m75G5290, m75G4552 and m75G4430, we have taken $m=7.5$ MeV and $G_{1}=5.290$, 4.552, 4.43 GeV-2, respectively. The (pseudo-)critical points are marked on the lines correspondingly.
In the framework of the NJL model, it is demonstrated that whether the first-order chiral phase transition occurs at $T=0$ (when $m\neq 0$) depends on the regularization scheme that is employed  (Buballa, 2005; Zhang et al., 2016; Kohyama et al., 2015). In Buballa (2005), the three-momentum cutoff regularization is used and a first-order phase transition happens at $T=0$. However, in Zhang et al. (2016), the authors use the PTR and find a crossover in the phase transition region at $T=0$. Actually, in Kohyama et al. (2015), it is clarified that the low current quark mass ($m\leq 4$ MeV) can result in a crossover at $T=0$ for both PTR and three-momentum cutoff regularization.
According to lattice QCD simulations, we notice that at zero temperature and under the PTR, when $m>7.5$ MeV, $\mu_{\rm c}$ of the modified NJL model is significantly lower than that in the normal NJL model. Sometimes even a first-order phase transition occurs. For example, from the phase diagram of the normal NJL model derived by Kohyama et al. (2015), we see that the crossover can only occur at $T=0$ when $m<10$ MeV. However, according to our calculations, a first-order phase transition has already occurred when $m>9$ MeV in the modified NJL model.
Figure 4: The number density of $u,d$ quarks versus chemical potentials $\mu$ at $T=0$. Line styles are the same as those in Fig. 3.
The dependence of quark number density on the chemical potential at $T=0$ is shown in Fig. 4. We can see that a larger $m$ leads to a lower quark number density when $T_{\rm pc}=181$ MeV and $\mu\in(220-290)$ MeV. For a fixed $m$ with different values of $G_{1}$, the quark number densities are only slightly different in the crossover region according to the lattice QCD simulation results on $T_{\rm pc}$. However, they are quite different in normal NJL models.
To describe the strongly interacting matter in hybrid stars, we need to consider the beta equilibrium and electric charge neutrality,
$\mu_{\rm d}=\mu_{\rm u}+\mu_{\rm e},$
$\frac{2}{3}\rho_{\rm u}-\frac{1}{3}\rho_{\rm d}-\rho_{\rm e}=0,$ (11)
where $\rho_{\rm e}=\mu_{\rm e}^{3}/3\pi^{2}$ is the number density of electrons at $T=0$. The relation between the baryon density $\rho_{\rm B}$ and baryon chemical potential $\mu_{\rm B}$ is shown in Fig. 5, where $\mu_{\rm B}=\mu_{\rm u}+2\mu_{\rm d}$ and $\rho_{\rm B}=(\rho_{\rm u}+\rho_{\rm d})/3$. In Fig. 5, the shape of each curve is quite similar to that in Fig. 4. Interestingly, in the modified NJL model, the $\rho_{\rm B}-\mu_{\rm B}$ relation of the quark system is almost identical even when different parameter sets are adopted. The difference mainly occurs in the phase transition region.
Figure 5: The baryon density versus baryon chemical potential of the quark system. Line styles are the same as those in Fig. 3.
According to the definition, the EOS of dense quark matter at $T=0$ is (Zong & Sun, 2008)
[…]
and the energy density of the quark system can be expressed as (Yan et al., 2012; Benvenuto & Lugones, 1995)
$\epsilon=-P+\sum_{i=u,d,e}\mu_{\rm i}\rho_{\rm i}.$ (13)
Note that $P(\mu=0)$ in Eq. (12) is irrelevant to the chemical potential. It represents the negative vacuum pressure, which corresponds to the vacuum bag constant ($-B$) in the MIT bag model. In the NJL model, it can be calculated by using Eq. (2.60) of Buballa (2005) under the mean field approximation. In this way, the bag constant is regarded as the pressure difference between the trivial and the non-trivial vacuum. It is known that the trivial and non-trivial vacuum can be described by the (quasi-)Wigner and Nambu solution of the gap equation, respectively. As suggested by Xu et al. (2018); Cui et al. (2018); Li et al. (2019a); Wang et al. (2019), the vacuum pressure is defined as
$P(\mu=0)$ $=$ $P(M_{N})-P(M_{W})$ (14)
$=$ $\Omega(0,0,\langle\bar{\psi}\psi\rangle_{W})-\Omega(0,0,\langle\bar{\psi}\psi\rangle_{N}),$
where $M_{N}$ and $M_{W}$ represents the Nambu and quasi-Wigner solution of the gap equation, respectively, and $\langle\bar{\psi}\psi\rangle_{N}$ and $\langle\bar{\psi}\psi\rangle_{W}$ are the corresponding quark condensates.
3 Structure of hybrid stars
Under the Maxwell construction scheme, the first-order hadron-quark phase transition occurs when the baryon chemical potentials and pressures of these two phases are equal,
$P_{\rm H}(\mu_{\rm B,c})=P_{\rm Q}(\mu_{\rm B,c}),$ (15)
where $\mu_{\rm B,c}$ is the critical baryon chemical potential of the hadron-quark phase transition, which is around 1.5 and 1.6 GeV in the modified NJL model and the normal NJL model, respectively. Note that the dense hadronic matter in this work is described by the APR EOS with $A18+\delta\nu+UIX^{\ast}$ interaction (Akmal et al., 1998), which is strongly favored by recent neutron star observations. The hybrid EOS can be written as
$P(\mu_{\rm B})=\left\{\begin{array}[]{lcl}P_{\rm H},\,\,\,when\,\,\mu_{\rm B}\leq\mu_{\rm B,c},\\ P_{\rm Q},\,\,\,when\,\,\mu_{\rm B}\geq\mu_{\rm B,c}.\end{array}\right.$ (16)
The corresponding energy density of the hybrid EOS is
$\epsilon(\mu_{\rm B})=\left\{\begin{array}[]{lcl}\epsilon_{\rm H},\,\,\,when\,\,\mu_{\rm B}\leq\mu_{\rm B,c},\\ \epsilon_{\rm Q},\,\,\,when\,\,\mu_{\rm B}\geq\mu_{\rm B,c},\end{array}\right.$ (17)
where $\epsilon_{\rm Q}$ is the energy density of the quark system.
In Sec. 2, the parameter sets have been preliminarily constrained by lattice simulations, and the results are shown in Table 1. Here they will be further constrained based on recent measurements on neutron star mass, radius, and tidal deformability. According to our calculations, a 2 $M_{\odot}$ neutron star can provide a strong constraint on the lower limit of $G_{1}$. It rules out the parameter sets with $m<5.8$ MeV and $m>9.5$ MeV. On one hand, when $m<5.8$ MeV, the most massive hybrid star mass cannot reach 2 $M_{\odot}$. On the other hand, when $m>9.5$ MeV, the pressure of quark matter is higher than that of hadronic matter, and the most massive quark star is also lighter than 2 $M_{\odot}$.
In the following, to study the influence of various parameters on the EOS and on the structure of hybrid stars, we choose five representative quark EOSs and their corresponding hybrid EOSs for comparison. The corresponding parameter sets are $m=5.8$ MeV with $G_{1}=3.485$ GeV-2, $m=7.5$ MeV with $G_{1}=4.430,4.552,5.290$ GeV-2, respectively, and $m=9.5$ MeV with $G_{1}=5.572$ GeV-2. Specifically, the parameter sets with $(m,G_{1})=$(5.8 MeV, 3.485 GeV-2), (7.5 MeV, 4.552 GeV-2), (9.5 MeV, 5.572 GeV-2) correspond to the case of $T_{\rm pc}=181$ MeV, from which we can see the influence of $m$ on the results. The parameter sets with $(m,G_{1})=$(7.5 MeV, 4.430 GeV-2), (7.5 MeV, 4.552 GeV-2), (7.5 MeV, 5.290 GeV-2) correspond to the case of $m=7.5$ MeV, from which we can find the influence of $G_{1}$ on the results.
According to the discussion on the vacuum pressure $P(\mu=0)$ in Eq. 14, the bag constant $B^{1/4}$ corresponding to the parameter sets of $(m,G_{1})$=(5.8 MeV, 3.485 GeV-2), (7.5 MeV, 4.430 GeV-2), (7.5 MeV, 4.552 GeV-2), (9.5 MeV, 5.572 GeV-2) can be calculated as 122.7 MeV, 120.9 MeV, 123.6 MeV, 125.4 MeV, respectively. Comparing with the normal NJL model with $(m,G_{1})$ = (7.5 MeV, 5.290 GeV-2) and $B^{1/4}=137.1$ MeV, the modified NJL model under the constraint of $T_{\rm pc}$ has a smaller $B^{1/4}$, which ranges in 120 – 125 MeV. For a fixed $T_{\rm pc}$, a larger $m$ will lead to a slightly larger $B^{1/4}$. Note that the range of $B$ in this work is consistent with that in  Song et al. (1992). Considering the medium effect, the study  (Lu et al., 1998) also indicated the decline of $B$ in the framework of the quark-meson coupling model.
In Fig. 6, the influences of $m$ and $G_{1}$ on the pressure of the quark matter are presented in the left panel and right panel with $T_{\rm pc}=181$ MeV and $m=7.5$ MeV, respectively. We can see that after considering the constraint of lattice simulation results on $T_{\rm pc}$, the EOSs of quark matter under different parameter sets are very close to each other. It could also be seen that the deconfinement phase transition occurs at about $\mu_{B}\sim 1.5$ GeV. The left panel shows the EOSs corresponding to $T_{\rm pc}=181$ MeV and $m=5.8$, 7.5, 9.5 MeV, respectively. Similarly, the right panel shows the EOSs corresponding to $m=7.5$ MeV and $G_{1}=5.290$, 4.552, 4.430 GeV-2, respectively. The EOSs with $G_{1}=4.430$ and 4.552 GeV-2 are very close to each other and their corresponding $T_{\rm pc}$ values are in accordance with the lattice simulation results. However, both of these two EOSs are quite different from that of $G_{1}=5.290$ GeV-2, which corresponds to the case of normal NJL model.
Figure 6: The pressure versus baryon chemical potential ($\mu_{\rm B}$) for $T_{\rm pc}=181$ MeV (left panel) and $m=7.5$ MeV (right panel), respectively. The pressure of the hadronic matter described by BSK21 is also presented for a comparison. Here, m58, m75, m95 refer to $m=5.8$, 7.5, 9.5 MeV, respectively, and G5290, G4552, G4430 refer to $G_{1}=5.290$, 4.552, 4.430 GeV-2, respectively.
In Fig. 7, we present the $\epsilon-P$ relations of the hadronic matter, the quark matter and hybrid EOSs with the Maxwell construction. Each point marked with “x” represents the critical point of the corresponding first-order phase transition, which is denoted as ($P_{\rm pt}$, $\epsilon_{\rm pt}$) hereafter. The other marked point on each hybrid EOS refers to the center of the most massive hybrid star. Here we denote it as ($P_{\rm c}$, $\epsilon_{\rm c}$). We can see that according to the constraint of lattice simulation results on $T_{\rm pc}$, $\epsilon_{\rm c}$ ($\epsilon_{\rm pt}$) of the hybrid EOSs only varies slightly among different parameter sets. Generally, a larger $G_{1}$ will lead to a larger value $P_{\rm c}$ and $\epsilon_{\rm c}$. The corresponding ($P_{\rm pt}$, $\epsilon_{\rm pt}$) and ($P_{\rm c}$, $\epsilon_{\rm c}$) points of five representative hybrid EOSs are listed in Table 2.
Table 2: The corresponding ($P_{\rm pt}$, $\epsilon_{\rm pt}$) and ($P_{\rm c}$, $\epsilon_{\rm c}$) points of five representative hybrid EOSs.
$\quad\,m$ $\quad\,\,G_{1}$ $\,\,\,\,\,(P_{pt},\epsilon_{pt})$ $\,\,\,\,\,\,\,(P_{\rm c},\epsilon_{\rm c})$
$\,\,[{\rm MeV}]$ $[{\rm GeV}^{-2}]$ $\,[{\rm MeV}\cdot{\rm fm}^{-3}]$ $\,[{\rm MeV}\cdot{\rm fm}^{-3}]$
   5.8     3.4853.485\,\,\,\,3.4853.485 $(299.9,856.5)$ $(348.6,1312.3)$
    4.4304.430\,\,\,\,4.4304.430 $(299.9,856.5)$ $(306.2,1179.6)$
   7.5     4.5524.552\,\,\,\,4.5524.552 $(299.9,856.5)$ $(332.5,1283.2)$
    5.2905.290\,\,\,\,5.2905.290 $(415.8,980.1)$ $(424.1,1707.7)$
   9.5     5.5725.572\,\,\,\,5.5725.572 $(326.6,886.3)$ $(350.5,1358.2)$
Figure 7: The $\epsilon-P$ relation of the quark matter and hybrid EOS when $T_{\rm pc}=181$ MeV (left panel) and $m=7.5$ MeV (right panel), respectively. The $\epsilon-P$ relation of the hadronic matter described by APR is shown by the red line. Here, m58, m75, m95 (APR-m58, APR-m75, APR-m95) refer to the quark matter (hybrid EOS) with $m=5.8$, 7.5, 9.5 MeV, respectively, and G5290, G4552, G4430 (APR-G5290, APR-G4552, APR-G4430) refer to quark matter (hybrid EOS) with $G_{1}=5.290$, 4.552, 4.430 GeV-2, respectively.
Once the EOS is determined, we can solve the Tolman-Oppenheimer-Volkoff (TOV) equation numerically to get the $M-R$ and mass-central energy density ($M-\epsilon_{\rm c}$) relations. In Fig. 8 and Fig. 9, to study the influence of $m$ and $G_{1}$ on the $M-R$ and $M-\epsilon_{\rm c}$ relations, the corresponding results for $T_{\rm pc}=181$ MeV and $m=7.5$ MeV are shown in the left panel and right panel, respectively. In Fig. 8, the most massive quark star for $m=5.8$, 7.5, 9.5 MeV ($G_{1}=5.290$, 4.552, 4.430 GeV-2) is about 1.892 $M_{\odot}$ (1.879 $M_{\odot}$), not satisfying the 2 $M_{\odot}$ constraint. In addition, the quark stars cannot fulfill the recent $M-R$ constraint from the NICER measurement of PSR J0030+0451 (Miller et al., 2019b), although the other constraint from PSR J0740+6620 (Miller et al., 2021) is marginally fulfilled by the quark EOSs with the parameter sets of $(m,G_{1})=(5.8,3.485)$ and $(7.5,4.430)$. However, the hybrid EOSs obtained with the Maxwell construction approach in this work can produce hybrid stars in consistent with these astronomical observations, although their quark matter cores are relatively small (about 0.03 $M_{\odot}$).
Figure 8: The $M-R$ relations of quark stars and hybrid stars for $T_{\rm pc}=181$ MeV (left panel) and $m=7.5$ MeV (right panel), respectively. Line styles are the same as in Fig. 7.
The $M-\epsilon_{\rm c}$ relations are shown in Fig. 9. We can find that for stable neutron stars (whether they are hadron stars, quark stars, or hybrid stars), a larger $\epsilon_{\rm c}$ corresponds to a more massive star. The $\epsilon_{\rm pt}$ ($\epsilon_{\rm c}$) of the hybrid stars whose corresponding hybrid EOSs satisfy the constraint of the lattice simulation results on $T_{\rm pc}$ is in a range of $856-886$ ($1180-1358$) MeV/fm3, which can also be seen in Table 2. However, for the normal NJL model with $m=7.5$ MeV, the corresponding $\epsilon_{\rm pt}$ ($\epsilon_{\rm c}$) of hybrid stars is about 980 (1700) MeV/fm3, quite different from the results of the modified NJL model.
Figure 9: The $M-\epsilon_{\rm c}$ relations of quark stars and hybrid stars for $T_{\rm pc}=181$ MeV (left panel) and $m=7.5$ MeV (right panel), respectively. Line styles are the same as in Fig. 7.
We have also calculated the tidal deformability of hadron stars, quark stars and hybrid stars in this work, which is defined as (Hinderer et al., 2010),
$\Lambda=\frac{2}{3}k_{2}R^{5}.$ (18)
Here $k_{2}$ is the dimensionless tidal Love number for $l=2$, which can be calculated by
$k_{2}$ $=\frac{8C^{5}}{5}(1-2C)^{2}[2+2C(y-1)-y]$ (19)
$\times$ $\{2C[6-3y+3C(5y-8)]$
$+$ $4C^{3}[13-11y+C(3y-2)+2C^{2}(1+y)]$
$+$ $3(1-2C)^{2}[2+2C(y-1)-y]ln(1-2C)\}^{-1},$
where $C=M/R$ refers to the compactness of the star, and $y$ is defined as
$y=R\beta(R)/H(R)-4\pi R^{3}\epsilon_{0}/M,$ (20)
where $\epsilon_{0}$ is the energy density at the surface of the star. The dimensionless parameter $y$ can be calculated by solving the differential equations  (Hinderer et al., 2010),
$\frac{dH}{dr}$ $=$ $\beta,$
$\frac{d\beta}{dr}$ $=$ $2(1-2\frac{m_{r}}{r})^{-1}H\{-2\pi[5\epsilon+9P+f(\epsilon+P)]$ (21)
$+$ $\frac{3}{r^{2}}+2(1-2\frac{m_{r}}{r})^{-1}(\frac{m_{r}}{r^{2}}+4\pi rP)^{2}\}$
$+$ $\frac{2\beta}{r}(1-2\frac{m_{r}}{r})^{-1}\{\frac{m_{r}}{r}+2\pi r^{2}(\epsilon-P)-1\},\,$
where $H(r)$ is the metric function, and $f={\rm d}\epsilon/{\rm d}P$.
The $\Lambda-M$ relation is shown in Fig. 10. We can see that for quark stars and hadron stars described by the modified NJL model and APR hadronic model, the corresponding values of $\Lambda(1.4M_{\odot})$ satisfy the constraint from GW170817, i.e., $\Lambda(1.4M_{\odot})=190_{-120}^{+390}$  (Abbott & et al., 2018). For stable hybrid stars whose maximum masses are higher than 2 $M_{\odot}$ in this work, the corresponding hybrid EOSs demonstrate that neutron stars with the masses lower than 1.974 $M_{\odot}$ are still hadron stars and the quark matter cores do not exist inside them. Therefore, the $\Lambda-M$ relations from these hybrid EOSs are the same with that of hadron stars when $M\leq 1.974$ $M_{\odot}$, and we do not show $\Lambda-M$ relations of hybrid stars in Fig. 10. In other words, according to the hybrid EOSs constrained in this work, the BNS in GW170817 whose masses are estimated to be $1.17-1.36$ and $1.36-1.60$ $M_{\odot}$ (Abbott & et al., 2017a), respectively, should both be hadron stars.
Figure 10: The $\Lambda-M$ relation of quark stars and hadron stars. Here, m58, m75, m95 refer to quark stars with $T_{\rm pc}=181$ MeV and $m=5.8$, 7.5, 9.5 MeV, respectively, and m75G5290, m75G4552, m75G4430 refer to quark stars with $m=7.5$ MeV and $G_{1}=5.290$, 4.552, 4.430 GeV-2, respectively. The observational constraint from GW170817, $\Lambda(1.4M_{\odot})=190_{-120}^{+390}$  (Abbott & et al., 2018) is also plotted for comparison.
For the sake of completeness, the $M-R$ properties of hybrid stars constructed by five representative hybrid EOSs are presented in Table 3, where ($R_{pt}$, $M_{pt}$) is related to the hadron-quark phase transition point, referring to the radius and mass of the most massive hadron star constructed by the hybrid EOS, and ($R_{\rm max}$, $M_{\rm max}$) is the radius and mass of the most massive hybrid star with a quark matter core. We can see that the masses of quark matter cores are in a range of $0.026-0.04$ $M_{\odot}$ in the framework of the modified NJL model. For a fixed $T_{\rm pc}$ ($m$), a larger $m$ ($G_{1}$) leads to a larger $M_{\rm max}$. Due to the constraint of $T_{\rm pc}$ from the lattice simulation results, $M_{\rm max}$ is about 0.1 $M_{\odot}$ lower than that of normal NJL model, while the quark matter core is about 0.015 $M_{\odot}$ heavier than that of normal NJL model.
Table 3: The corresponding ($R_{pt}$, $M_{pt}$) and ($R_{\rm max}$, $M_{\rm max}$) points of five representative hybrid EOSs.
$\quad\,m$ $\quad\,\,G_{1}$ $\,\,\,\,(R_{pt},M_{pt})$ $(R_{\rm max},M_{\rm max})$
$\,\,[{\rm MeV}]$ $[{\rm GeV}^{-2}]$ $\,\,\,\,\,[km,M_{\odot}]$ $\,\,\,\,\,\,[km,M_{\odot}]$
   5.8     3.4853.485\,\,\,\,3.4853.485 $(11.11,1.974)$ $\,(10.99,2.003)$
    4.4304.430\,\,\,\,4.4304.430 $(11.11,1.974)$ $\,(11.07,2.005)$
   7.5     4.5524.552\,\,\,\,4.5524.552 $(11.11,1.974)$ $\,(11.02,2.014)$
    5.2905.290\,\,\,\,5.2905.290 $(10.89,2.084)$ $\,(10.85,2.103)$
   9.5     5.5725.572\,\,\,\,5.5725.572 $(11.05,2.006)$ $\,(10.99,2.032)$
4 Summary
In this study, the modified 2-flavor NJL model and the APR EOS with $A18+\delta\nu+UIX^{\ast}$ interaction are introduced to investigate the nonstrange quark matter and hadronic matter in hybrid stars in light of a hypothesis that the quark matter may not be strange. To construct hybrid EOSs, the first-order hadron-quark phase transition and the corresponding Maxwell construction are considered.
It is noted that the modification of the coupling constant $G$ in the normal NJL model is helpful, because it is not only consistent with the QCD requirement in essence, but also in agreement with the lattice simulation results of $T_{\rm c}$. In the 2-flavor case, when $T_{\rm c}=173\pm 8$ MeV, the parameter space of $G_{1}$ can be constrained for a given current quark mass $m$, and the corresponding range of $G_{2}$ can also be determined. The results are shown in Table 1, implying that the normal 2-flavor NJL model with the four-quark scalar interaction (corresponding to the case of $G_{2}=0$ in our modified NJL model) is inconsistent with the lattice simulation results. For hybrid EOSs, the influence of $m$ and $G_{1}$ is very small when the constraint of $T_{\rm c}=173\pm 8$ MeV is taken into account. The hybrid EOSs derived from the modified NJL model are quite different from that of normal NJL model.
Considering astronomical observations and the stability of hybrid stars, the parameter $m$ is constrained to be $5.8-9.5$ MeV, and the parameter space of $G_{1}$ can also be determined. The quark EOSs constructed with the modified NJL model in this work is soft, and thus cannot satisfy the 2 $M_{\odot}$ constraint of neutron stars and some $M-R$ constraints from NICER missions. It is noted that in some previous studies, the quark matter cores may not exist in compact stars under the Maxwell construction  (Özel, 2006; Hoyos et al., 2016; Qin et al., 2023), or the maximum mass of quark matter cores may be larger than 0.6 $M_{\odot}$ and the BNS in GW170817 can be hybrid stars  (Ayriyan et al., 2021; Blaschke et al., 2020; Ayriyan et al., 2019; Pfaff et al., 2022). However, with the modified 2-flavor NJL model in this work, the hybrid EOSs with first-order hadron-quark transitions are still in agreement with current neutron star astronomical observations, and pure nonstrange quark matter cores can exist in hybrid stars, possessing a relatively small mass of $0.026-0.04$ $M_{\odot}$. Due to the constraint of $T_{\rm c}=173\pm 8$ MeV, $M_{\rm max}$ of hybrid stars will be about 0.1 $M_{\odot}$ lower than that of normal NJL model. According to the hybrid EOSs constrained in this work, the BNS in GW170817 whose masses are estimated to be $1.17-1.36$ and $1.36-1.60$ $M_{\odot}$ (Abbott & et al., 2017a), respectively, may be hadron stars.
We thank the anonymous referee for useful suggestions that led to an overall improvement of the presentation. This study is supported in part by the National Key Program for Science and Technology Research Development (2023YFB3002500), the National Natural Science Foundation of China (under Grants No. 12005192, No. 12075213, and No. 12233002), the Project funded by China Postdoctoral Science Foundation (No. 2020M672255, No. 2020TQ0287), National SKA Program of China No. 2020SKA0120300, the National Key R&\&&D Program of China (2021YFA0718500), the Natural Science Foundation of Henan Province of China (No. 242300421375), the Natural Science Foundation for Distinguished Young Scholars of Henan Province under grant number 242300421046, the start-up funding from Zhengzhou University. Y.F.H also acknowledges the support from the Xinjiang Tianchi Program.